# Andreas Schlaginhaufen SYCAMORE, EPFL &Reda Ouhamma SYCAMORE, EPFL &Maryam Kamgarpour

### v1

@@ A.4 Miscellaneous
$\Delta_{t}(\theta^{*})\mathbbm{1}(E_{t})\leq\mathbb{E}\left[{\Delta_{t}(\tilde{\theta}_{t})\mathbbm{1}(E_{t})\mid\mathcal{F}_{t},O_{t}}\right].$
Putting this together, we get
$r_{t}^{\text{TS}}\mathbbm{1}(E_{t})$ $\leq\mathbb{E}\left[{\left({\Delta_{t}(\tilde{\theta}_{t})-\underline{\Delta}_{t}}\right)\mathbbm{1}(E_{t})\mid\mathcal{F}_{t},O_{t}}\right]$
$\stackrel{{\scriptstyle(i)}}{{\leq}}\mathbb{E}\left[{\left({\langle\tilde{\theta}_{t},\phi(\pi_{t})-\phi(\pi_{t}^{\prime})\rangle+\varepsilon-\max_{\pi}\langle\underline{\theta}_{t},\phi(\pi)-\phi(\pi_{t}^{\prime})\rangle}\right)\mathbbm{1}(E_{t})\mid\mathcal{F}_{t},O_{t}}\right]$
=$\stackrel{{\scriptstyle(ii)}}{{=}}\mathbb{E}\left[{\left({\langle\tilde{\theta}_{t},\phi(\pi_{t})-\phi(\pi_{t}^{\prime})\rangle-\langle\underline{\theta}_{t},\phi(\pi_{t})-\phi(\pi_{t}^{\prime})\rangle}\right)\mathbbm{1}(E_{t})\mid\mathcal{F}_{t},O_{t}}\right]+\varepsilon$
$\stackrel{{\scriptstyle(iii)}}{{\leq}}2\beta_{t}(\delta^{\prime})c(\delta^{\prime\prime})\mathbb{E}\left[{\left\lVert\phi(\pi_{t})-\phi(\pi^{\prime}_{t})\right\rVert_{V_{t}^{-1}}\mid\mathcal{F}_{t},\widehat{E}_{t},O_{t}}\right]\Pr\left[{\widehat{E}_{t}\mid\mathcal{F}_{t}}\right]+\varepsilon,$
@@
$\mathbb{E}\left[{\left\lVert\phi(\pi_{t})-\phi(\pi^{\prime}_{t})\right\rVert_{V_{t}^{-1}}\mid\mathcal{F}_{t},\widehat{E}_{t},O_{t}}\right]\leq\mathbb{E}\left[{\left\lVert\phi(\pi_{t})-\phi(\pi^{\prime}_{t})\right\rVert_{V_{t}^{-1}}\mid\mathcal{F}_{t},\widehat{E}_{t}}\right]/\Pr\left[{O_{t}|\mathcal{F}_{t},\widehat{E}_{t}}\right].$
@@
$\Pr\left[{O_{t}\mid\mathcal{F}_{t},\widehat{E}_{t}}\right]\geq\Pr\left[{\Delta_{t}(\tilde{\theta}_{t})\geq\max_{\theta\in\mathcal{E}_{t}}\Delta_{t}(\theta)\mid\mathcal{F}_{t}}\right].$
Since $\theta\mapsto\Delta_{t}(\theta)$ is the sum of a linear function and the function $\theta\mapsto\max_{\pi}\langle\theta,\phi(\pi)\rangle$ which is convex and continuous by Proposition A.7. Then, applying Lemma A.3 with $f(\theta)=\Delta_{t}(\theta)$, $\tilde{x}=\tilde{\theta}_{t}\mid\mathcal{F}_{t}$, and $\mathcal{E}=\mathcal{E}_{t}$, we have
$\Pr\left[{O_{t}\mid\mathcal{F}_{t},\widehat{E}_{t}}\right]\geq 1/\left({4\sqrt{e\pi}}\right)=:p.$
@@
$r_{t}^{\text{TS}}\mathbbm{1}(E_{t})$ $\leq\frac{2\beta_{t}(\delta^{\prime})c(\delta^{\prime\prime})\mathbb{E}\left[{\left\lVert\phi(\pi_{t})-\phi(\pi^{\prime}_{t})\right\rVert_{V_{t}^{-1}}\mid\mathcal{F}_{t}}\right]}{p}+\varepsilon$
$\leq\frac{2\beta_{t}(\delta^{\prime})c(\delta^{\prime\prime})\mathbb{E}\left[{\left\lVert\phi(\tau_{t})-\phi(\tau^{\prime}_{t})\right\rVert_{V_{t}^{-1}}\mid\mathcal{F}_{t}}\right]}{p}+\varepsilon,$
where the second inequality follows from the convexity of norms. Applying Lemma A.5 2), we have on the good event $G$ that
$\sum_{t=1}^{T}r_{t}^{\text{TS}}$ $\leq\dfrac{2\beta_{T}(\delta^{\prime})c(\delta^{\prime\prime})}{p}\sum_{t=1}^{T}\mathbb{E}\left[{\left\lVert\phi(\tau_{t})-\phi(\tau^{\prime}_{t})\right\rVert_{V_{t}^{-1}}\mid\mathcal{F}_{t}}\right]+T\varepsilon$ (5)
@@
$\leq\beta_{t}(\delta^{\prime})(1+c(\delta^{\prime\prime}))\mathbb{E}\left[{\left\lVert\phi(\tau_{t})-\phi(\tau^{\prime}_{t})\right\rVert_{V_{t}^{-1}}|\mathcal{F}_{t}^{\theta}}\right],$
where the last inequality follows from the convexity of norms. From here, we can again apply the second result of Lemma A.5. We deduce that on the good event $G$, we have
$\sum_{t=1}^{T}r_{t}^{\text{MLE}}$ $\leq\beta_{T}(\delta^{\prime})(1+c(\delta^{\prime\prime}))\sum_{t=1}^{T}\mathbb{E}\left[{\left\lVert\phi(\tau_{t})-\phi(\tau^{\prime}_{t})\right\rVert_{V_{t}^{-1}}|\mathcal{F}_{t}^{\theta}}\right]$
@@
Proof.
Recall that in Algorithm 1, we set $\varepsilon=1/\sqrt{T}$, $\delta^{\prime}=\delta/5$, and $\hat{\pi}=\texttt{A}^{\texttt{PAC}}_{\texttt{RL}}(\bar{\theta}_{T},\varepsilon,\delta^{\prime})$ for $\bar{\theta}_{T}=\frac{1}{T}\sum_{t=1}^{T}\hat{\theta}_{t}$. Moreover, as $\alpha=0$, we further have $\tilde{\theta}_{t}\mid\mathcal{F}_{t}\sim\mathcal{N}(0,\beta_{t}^{2}V_{t}^{-1})$, and choose the policies $\pi_{t}=\texttt{A}^{\texttt{PAC}}_{\texttt{RL}}(\tilde{\theta}_{t})$ and $\pi_{t}^{\prime}=\pi_{t-1}$.
In the following, we will denote $\pi^{*}\in\operatorname*{arg\,max}_{\pi}\langle\theta^{*},\phi(\pi)\rangle$ for an arbitrary optimal policy corresponding to the ground truth reward parameter $\theta^{*}$. Under the good event $G$ of Proposition B.1, we can bound the suboptimality of $\hat{\pi}$ as follows
$\mathsf{SubOpt}(\hat{\pi})$ $=\langle\theta^{*},\phi(\pi^{*})-\phi(\hat{\pi})\rangle$
$\leq\langle\theta^{*}-\bar{\theta}_{T},\phi(\pi^{*})-\phi(\hat{\pi})\rangle+\varepsilon$
$=\varepsilon+\frac{1}{T}\sum_{t=1}^{T}\langle\theta^{*}-\hat{\theta}_{t},\phi(\pi^{*})-\phi(\hat{\pi})\rangle,$
where we used $\varepsilon$-optimality of $\hat{\pi}$ and the definition of $\bar{\theta}_{T}$. Now, recall that conditioned on $E_{t}$, we have $\theta^{*}\in\mathcal{E}_{t}$ and $\hat{\theta}_{t}+\tilde{\theta}_{t}\in{\mathcal{E}}_{t}^{\text{TS}}$. Define ¯$\bar{\mathcal{E}}_{t}=\{\theta:\theta+\hat{\theta}\in\mathcal{E}_{t}\}$. Then, on the good event $G$ it holds that:
$\langle\theta^{*}-\hat{\theta}_{t},\phi(\pi^{*})-\phi(\hat{\pi})\rangle\mathbbm{1}\left({E_{t}}\right)$ $\leq\max_{\pi,\pi^{\prime}}\max_{\theta\in\mathcal{E}_{t}}\langle\theta-\hat{\theta}_{t},\phi(\pi)-\phi(\pi^{\prime})\rangle\mathbbm{1}\left({\widetilde{E}_{t}}\right)$
$=\max_{\pi,\pi^{\prime}}\max_{\theta\in\bar{\mathcal{E}}_{t}}\langle\theta,\phi(\pi)-\phi(\pi^{\prime})\rangle\mathbbm{1}\left({\widetilde{E}_{t}}\right)$
=$\stackrel{{\scriptstyle(i)}}{{=}}\beta_{t}\max_{\pi,\pi^{\prime}}\left\lVert\phi(\pi)-\phi(\pi^{\prime})\right\rVert_{V_{t}^{-1}}\mathbbm{1}\left({\widetilde{E}_{t}}\right)$
$\stackrel{{\scriptstyle(ii)}}{{\leq}}\beta_{t}\max_{\pi,\pi^{\prime}}\left({\left\lVert\phi(\pi)-\phi(\pi_{t}^{\prime})\right\rVert_{V_{t}^{-1}}+\left\lVert\phi(\pi_{t}^{\prime})-\phi(\pi^{\prime})\right\rVert_{V_{t}^{-1}}}\right)\mathbbm{1}\left({\widetilde{E}_{t}}\right)$
$=2\beta_{t}\max_{\pi}\left\lVert\phi(\pi)-\phi(\pi_{t}^{\prime})\right\rVert_{V_{t}^{-1}}\mathbbm{1}\left({\widetilde{E}_{t}}\right)$
=$\stackrel{{\scriptstyle(iii)}}{{=}}2\max_{\pi}\max_{\theta\in\bar{\mathcal{E}}_{t}}\langle\theta,\phi(\pi)-\phi(\pi_{t}^{\prime})\rangle\mathbbm{1}\left({\widetilde{E}_{t}}\right)$
$=2\max_{\theta\in\bar{\mathcal{E}}_{t}}\max_{\pi}\langle\theta,\phi(\pi)-\phi(\pi_{t}^{\prime})\rangle\mathbbm{1}\left({\widetilde{E}_{t}}\right).$
Here, $(i)$ and $(iii)$ follow because ¯$\bar{\mathcal{E}}_{t}$ is a centred ellipsoid and $(ii)$ from the triangle inequality. Next, we define the non-negative function $f_{t}(\theta):=\max_{\pi}\langle\theta,\phi(\pi)-\phi(\pi_{t}^{\prime})\rangle$ and the event $O_{t}=\{f_{t}(\tilde{\theta}_{t})\geq\max_{\theta\in\bar{\mathcal{E}}_{t}}f_{t}(\theta)\}$. By Proposition A.7 $f_{t}$ is convex and continuous. Applying Lemma A.3 with $f=f_{t}$, $\tilde{x}=\tilde{\theta}_{t}\mid\mathcal{F}_{t}$, and =¯$\mathcal{E}=\bar{\mathcal{E}}_{t}$, it holds that $\Pr\left[{O_{t}\mid\mathcal{F}_{t}}\right]\geq p:=1/(4\sqrt{e\pi})$. Hence,
$\langle\theta^{*}-\hat{\theta}_{t},\phi(\pi^{*})-\phi(\hat{\pi})\rangle\mathbbm{1}\left({E_{t}}\right)$ $\leq 2\max_{\theta\in\bar{\mathcal{E}}_{t}}\max_{\pi}\langle\theta,\phi(\pi)-\phi(\pi_{t}^{\prime})\rangle\mathbbm{1}\left({\widetilde{E}_{t}}\right)$
$\leq 2\mathbb{E}\left[{\max_{\pi}\langle\tilde{\theta}_{t},\phi(\pi)-\phi(\pi_{t}^{\prime})\rangle\mathbbm{1}\left({\widetilde{E}_{t}}\right)\mid O_{t},\mathcal{F}_{t}}\right]$
$\leq 2\mathbb{E}\left[{\langle\tilde{\theta}_{t},\phi(\pi_{t})-\phi(\pi_{t}^{\prime})\rangle\mathbbm{1}\left({\widetilde{E}_{t}}\right)\mid O_{t},\mathcal{F}_{t}}\right]+2\varepsilon$
$\leq 2\beta_{t}(\delta^{\prime})c(\delta^{\prime\prime})\mathbb{E}\left[{\left\lVert\phi(\pi_{t})-\phi(\pi_{t}^{\prime})\right\rVert_{V_{t}^{-1}}\mid O_{t},\mathcal{F}_{t}}\right]+2\varepsilon$
$\leq\frac{2\beta_{t}(\delta^{\prime})c(\delta^{\prime\prime})}{p}\mathbb{E}\left[{\left\lVert\phi(\pi_{t})-\phi(\pi_{t}^{\prime})\right\rVert_{V_{t}^{-1}}\mid\mathcal{F}_{t}}\right]+2\varepsilon.$
Using Lemma A.5, we conclude that with probability $1-\delta$, we have that
$\mathsf{SubOpt}(\hat{\pi})\leq 3\varepsilon+\frac{2\beta_{t}(\delta^{\prime})c(\delta^{\prime\prime})}{pT}\sum_{t=1}^{T}\mathbb{E}\left[{\left\lVert\phi(\pi_{t})-\phi(\pi_{t}^{\prime})\right\rVert_{V_{t}^{-1}}\mid\mathcal{F}_{t}}\right]$
$\leq$ $\frac{3}{\sqrt{T}}+\dfrac{2\left[{\sqrt{\kappa\left[{\log\left({\frac{5}{\delta}}\right)+d\log\left({\max\left\{{e,\frac{4eBLH_{\gamma}(T-1)}{d}}\right\}}\right)}\right]}+2\sqrt{\lambda}B}\right]\sqrt{2d\log(10dT/\delta)}}{pT}$
Appendix D Proof of Theorem 4.1
@@
Lemma D.1 (Number of design matrix updates).
@@
$\sum_{t=1}^{T}\mathbbm{1}\{V_{t}\neq V_{t-1}\}\leq\frac{d}{\log\left({\frac{1+C}{1-1/e}}\right)}\log\left({1+\frac{T(LH_{\gamma})^{2}}{\lambda d}}\right).$

### v2

@@ Acknowledgments
$\Delta_{t}(\theta^{*})\leq\mathbb{E}\left[{\Delta_{t}(\tilde{\theta}_{t})\mid\mathcal{F}_{t-1},O_{t}}\right].$
Putting this together, while keeping track of the events $E_{t}$ and $A_{t}$, we have
$r_{t}^{\text{TS}}\mathbbm{1}(E_{t}\cap A_{t})$ $\leq\mathbb{E}\left[{\left({\Delta_{t}(\tilde{\theta}_{t})-\underline{\Delta}_{t}}\right)\mathbbm{1}(E_{t}\cap A_{t})\mid\mathcal{F}_{t-1},O_{t}}\right]$
$\stackrel{{\scriptstyle(i)}}{{\leq}}\mathbb{E}\left[{\left({\langle\tilde{\theta}_{t},\phi(\pi_{t})-\phi(\pi_{t}^{\prime})\rangle+\varepsilon-\max_{\pi}\langle\underline{\theta}_{t},\phi(\pi)-\phi(\pi_{t}^{\prime})\rangle}\right)\mathbbm{1}(E_{t}\cap A_{t})\mid\mathcal{F}_{t-1},O_{t}}\right]$
$\stackrel{{\scriptstyle(ii)}}{{\leq}}\mathbb{E}\left[{\left({\langle\tilde{\theta}_{t},\phi(\pi_{t})-\phi(\pi_{t}^{\prime})\rangle-\langle\underline{\theta}_{t},\phi(\pi_{t})-\phi(\pi_{t}^{\prime})\rangle}\right)\mathbbm{1}(E_{t}\cap A_{t})\mid\mathcal{F}_{t-1},O_{t}}\right]+\varepsilon$
$\stackrel{{\scriptstyle(iii)}}{{\leq}}2\beta_{t}(\delta^{\prime})c(\delta^{\prime\prime})\mathbb{E}\left[{\left\lVert\phi(\pi_{t})-\phi(\pi^{\prime}_{t})\right\rVert_{V_{t}^{-1}}\mid\mathcal{F}_{t-1},\widehat{E}_{t},O_{t}}\right]\Pr\left[{\widehat{E}_{t}\mid\mathcal{F}_{t-1}}\right]+\varepsilon,$
@@
$\mathbb{E}\left[{\left\lVert\phi(\pi_{t})-\phi(\pi^{\prime}_{t})\right\rVert_{V_{t}^{-1}}\mid\mathcal{F}_{t-1},\widehat{E}_{t},O_{t}}\right]\leq\mathbb{E}\left[{\left\lVert\phi(\pi_{t})-\phi(\pi^{\prime}_{t})\right\rVert_{V_{t}^{-1}}\mid\mathcal{F}_{t-1},\widehat{E}_{t}}\right]/\Pr\left[{O_{t}|\mathcal{F}_{t-1},\widehat{E}_{t}}\right].$
@@
$\Pr\left[{O_{t}\mid\mathcal{F}_{t-1},\widehat{E}_{t}}\right]\geq\Pr\left[{\Delta_{t}(\tilde{\theta}_{t})\geq\max_{\theta\in\mathcal{E}_{t}}\Delta_{t}(\theta)\mid\mathcal{F}_{t-1}}\right].$
Since $\theta\mapsto\Delta_{t}(\theta)$ is the sum of a linear function and the function $\theta\mapsto\max_{\pi}\langle\theta,\phi(\pi)\rangle$ which is convex and continuous by Proposition D.7. Then, applying Lemma D.2 with $f(\theta)=\Delta_{t}(\theta)$, $\tilde{x}=\tilde{\theta}_{t}\mid\mathcal{F}_{t-1}$, and $\mathcal{E}=\mathcal{E}_{t}$, we have
$\Pr\left[{O_{t}\mid\mathcal{F}_{t-1},\widehat{E}_{t}}\right]\geq 1/\left({4\sqrt{e\pi}}\right)=:p.$
@@
$r_{t}^{\text{TS}}\mathbbm{1}(E_{t}\cap A_{t})$ $\leq\frac{2\beta_{t}(\delta^{\prime})c(\delta^{\prime\prime})\mathbb{E}\left[{\left\lVert\phi(\pi_{t})-\phi(\pi^{\prime}_{t})\right\rVert_{V_{t}^{-1}}\mid\mathcal{F}_{t-1}}\right]}{p}+\varepsilon$
$\leq\frac{2\beta_{t}(\delta^{\prime})c(\delta^{\prime\prime})\mathbb{E}\left[{\left\lVert\phi(\tau_{t})-\phi(\tau^{\prime}_{t})\right\rVert_{V_{t}^{-1}}\mid\mathcal{F}_{t-1}}\right]}{p}+\varepsilon,$
where the second inequality follows from the convexity of norms. Applying Lemma D.4 2), we have on the good event $G$ that
$\sum_{t=1}^{T}r_{t}^{\text{TS}}$ $\leq\dfrac{2\beta_{T}(\delta^{\prime})c(\delta^{\prime\prime})}{p}\sum_{t=1}^{T}\mathbb{E}\left[{\left\lVert\phi(\tau_{t})-\phi(\tau^{\prime}_{t})\right\rVert_{V_{t}^{-1}}\mid\mathcal{F}_{t-1}}\right]+T\varepsilon$ (3)
@@
$\leq\beta_{t}(\delta^{\prime})(1+c(\delta^{\prime\prime}))\mathbb{E}\left[{\left\lVert\phi(\tau_{t})-\phi(\tau^{\prime}_{t})\right\rVert_{V_{t}^{-1}}|\mathcal{F}_{t-1}^{\theta}}\right],$
where the last inequality follows from the convexity of norms. From here, we can again apply the second result of Lemma D.4. We deduce that on the good event $G$, we have
$\sum_{t=1}^{T}r_{t}^{\text{MLE}}$ $\leq\beta_{T}(\delta^{\prime})(1+c(\delta^{\prime\prime}))\sum_{t=1}^{T}\mathbb{E}\left[{\left\lVert\phi(\tau_{t})-\phi(\tau^{\prime}_{t})\right\rVert_{V_{t}^{-1}}|\mathcal{F}_{t-1}^{\theta}}\right]$
@@
Appendix B Proof of Theorem 3.2.2
[…]
Proof.
Similar to the proof of Theorem 4.2, we start by defining a set of good events.
Step 1 (good events): Recall that in Algorithm 2 we set $\varepsilon=1/\sqrt{T}$ and $\delta^{\prime}=\delta/5$; at round $t$, conditional on $\mathcal{F}_{t-1}=\sigma(x_{1},\ldots,x_{t-1})$, we sample $\tilde{\theta}_{t}\sim\mathcal{N}(0,V_{t}^{-1})$ with $V_{t}=\lambda I+\sum_{k=1}^{t-1}x_{k}x_{k}^{\top}$ (where $x_{k}=\phi(\tau_{k})-\phi(\tau_{k}^{\prime})$), and $\hat{\theta}$ is the MLE estimate at time $T+1$ using all the data from round $1$ to $T$. We now define the following high probability events.
  1. 1.
    Let $\delta^{\prime\prime}=\delta^{\prime}/T$ and $c(\delta^{\prime\prime}):=\sqrt{2d\log(2d/\delta^{\prime\prime})}=\sqrt{2d\log(10dT/\delta)}$. Consider the centered ellipsoid
    $\bar{\mathcal{E}}_{t}^{\text{TS}}:=\left\{{\theta\in\mathbb{R}^{d}:\left\lVert\theta\right\rVert_{V_{t}}\leq c(\delta^{\prime\prime})}\right\}.$
    We define the events
    $\widehat{E}:=\left\{{\|\theta^{*}-\hat{\theta}\|_{V_{T+1}}\leq\beta_{T+1}(\delta^{\prime})}\right\},\quad\widetilde{E}_{t}:=\left\{{\tilde{\theta}_{t}\in\bar{\mathcal{E}}_{t}^{\text{TS}}}\right\},$
    and let $G_{1}:=\bigcap_{t=1}^{T}\widetilde{E}_{t}\cap\widehat{E}$. This event implies that $\theta^{*}$ lies in the confidence ellipsoid $\mathcal{E}_{T+1}(\delta^{\prime})$, and the sampled parameter lies in $\bar{\mathcal{E}}_{t}^{\text{TS}}$ for all $t\in[1,T]$.
  2. 2.
    Let $G_{2}$ denote the event that for
    it holds that
    $\sum_{t=1}^{T}\mathbb{E}\left[{\left\lVert\phi(\tau_{t})-\phi(\tau_{t}^{\prime})\right\rVert_{V_{t}^{-1}}\mid\mathcal{F}_{t-1}}\right]\leq C_{T}.$
  3. 3.
    Let $A_{t}=\left\{{V_{\tilde{\theta}_{t}}^{*}-V_{\tilde{\theta}_{t}}^{\pi_{t}}\leq\varepsilon}\right\}$, where $\pi_{t}=\texttt{A}^{\texttt{PAC}}_{\texttt{RL}}\left({\tilde{\theta}_{t},\varepsilon,\delta^{\prime}/T}\right)$, and let $G_{3}=\bigcap_{t=1}^{T}A_{t}$.
  4. 4.
    Let $G_{4}$ denote the event that $V_{\hat{\theta}}^{*}-V_{\hat{\theta}}^{\hat{\pi}}\leq\varepsilon$ where $\hat{\pi}=\texttt{A}^{\texttt{PAC}}_{\texttt{RL}}\left({\hat{\theta},\varepsilon,\delta^{\prime}}\right)$.
Analogously to Proposition A.1, it follows that $\Pr[G_{1}^{\complement}]\leq 2\delta^{\prime}$ and $\Pr[G_{k}^{\complement}]\leq\delta^{\prime}$ for $k=2,3,4$. Thus, by a union bound, the good event $G:=\bigcap_{i=1}^{4}G_{i}$ happens with probability at least $1-\delta$.
Step 2 (suboptimality bound): In the following, we will denote $\pi^{*}\in\operatorname*{arg\,max}_{\pi}\langle\theta^{*},\phi(\pi)\rangle$ for an arbitrary optimal policy corresponding to the ground truth reward parameter $\theta^{*}$. Under the above good event $G$, we can bound the suboptimality of $\hat{\pi}$ as follows
[…]
$\leq\max_{\pi,\pi^{\prime}}\langle\theta^{*}-\hat{\theta},\phi(\pi)-\phi(\pi^{\prime})\rangle+\varepsilon\leq\beta_{T+1}(\delta^{\prime})\max_{\pi,\pi^{\prime}}\left\lVert\phi(\pi)-\phi(\pi^{\prime})\right\rVert_{V_{T+1}^{-1}}+\varepsilon.$ (4)
Next, we continue with bounding the term $\max_{\pi,\pi^{\prime}}\left\lVert\phi(\pi)-\phi(\pi^{\prime})\right\rVert_{V_{T+1}^{-1}}$.
Step 3 (approximate G-optimal design): Due to the matrix inequalities $V_{T+1}\succeq V_{T}\succeq\ldots\succeq V_{1}$, we have
$\left\lVert\cdot\right\rVert_{V_{T+1}^{-1}}\leq\left\lVert\cdot\right\rVert_{V_{T}^{-1}}\leq\ldots\leq\left\lVert\cdot\right\rVert_{V_{1}^{-1}}.$
Therefore, we can bound the right-hand-side of (4) as follows
$\max_{\pi,\pi^{\prime}}\left\lVert\phi(\pi)-\phi(\pi^{\prime})\right\rVert_{V_{T+1}^{-1}}$ $\leq\frac{1}{T}\sum_{t=1}^{T}\max_{\pi,\pi^{\prime}}\left\lVert\phi(\pi)-\phi(\pi^{\prime})\right\rVert_{V_{T+1}^{-1}}$
$\leq\frac{1}{T}\sum_{t=1}^{T}\max_{\pi,\pi^{\prime}}\left\lVert\phi(\pi)-\phi(\pi^{\prime})\right\rVert_{V_{t}^{-1}}$
$\leq\frac{2}{T}\sum_{t=1}^{T}\max_{\pi}\left\lVert\phi(\pi)-\phi(\pi_{t}^{\prime})\right\rVert_{V_{t}^{-1}},$
where we used the triangle inequality in the last step. If it were tractable to compute $\pi_{t}=\operatorname*{arg\,max}_{\pi}\left\lVert\phi(\pi)-\phi(\pi_{t}^{\prime})\right\rVert_{V_{t}^{-1}}$, we could invoke Lemma D.4 directly from here. Instead, we leverage randomized exploration to keep our algorithm tractable. In particular, we consider the function $f_{t}(\theta)=\max_{\pi}\langle\theta,\phi(\pi)-\phi(\pi_{t}^{\prime})\rangle$, the ellipsoid $\bar{\mathcal{E}}_{t}=\left\{{\theta:\left\lVert\theta\right\rVert_{V_{t}}\leq 1}\right\}$, and the event $O_{t}=\{f_{t}(\tilde{\theta}_{t})\geq\max_{\theta\in\bar{\mathcal{E}}_{t}}f_{t}(\theta)\}$. By Proposition D.7 $f_{t}$ is convex and continuous. Applying Lemma D.2 with $f=f_{t}$, $\tilde{\theta}=\tilde{\theta}_{t}\mid\mathcal{F}_{t-1}$, and $\mathcal{E}=\bar{\mathcal{E}}_{t}$, it holds that $\Pr\left[{O_{t}\mid\mathcal{F}_{t-1}}\right]\geq p:=1/(4\sqrt{e\pi})$. Hence, we can proceed analogously to the proof of Theorem 3.2.1:
$\max_{\pi}\left\lVert\phi(\pi)-\phi(\pi_{t}^{\prime})\right\rVert_{V_{t}^{-1}}\mathbbm{1}\left({\widetilde{E}_{t}\cap A_{t}}\right)$ $\stackrel{{\scriptstyle(i)}}{{=}}\max_{\theta\in\bar{\mathcal{E}}_{t}}f_{t}(\theta)\mathbbm{1}\left({\widetilde{E}_{t}\cap A_{t}}\right)$
$\leq\mathbb{E}\left[{f_{t}(\tilde{\theta}_{t})\mathbbm{1}\left({\widetilde{E}_{t}\cap A_{t}}\right)\mid O_{t},\mathcal{F}_{t-1}}\right]$
$\leq\mathbb{E}\left[{\langle\tilde{\theta}_{t},\phi(\pi_{t})-\phi(\pi_{t}^{\prime})\rangle\mathbbm{1}\left({\widetilde{E}_{t}\cap A_{t}}\right)\mid O_{t},\mathcal{F}_{t-1}}\right]+\varepsilon$
$\leq c(\delta^{\prime\prime})\mathbb{E}\left[{\left\lVert\phi(\pi_{t})-\phi(\pi_{t}^{\prime})\right\rVert_{V_{t}^{-1}}\mid O_{t},\mathcal{F}_{t-1}}\right]+\varepsilon$
$\stackrel{{\scriptstyle(ii)}}{{\leq}}\frac{c(\delta^{\prime\prime})}{p}\mathbb{E}\left[{\left\lVert\phi(\pi_{t})-\phi(\pi_{t}^{\prime})\right\rVert_{V_{t}^{-1}}\mid\mathcal{F}_{t-1}}\right]+\varepsilon$
$\leq\frac{c(\delta^{\prime\prime})}{p}\mathbb{E}\left[{\left\lVert\phi(\tau_{t})-\phi(\tau_{t}^{\prime})\right\rVert_{V_{t}^{-1}}\mid\mathcal{F}_{t-1}}\right]+\varepsilon.$
Here, $(i)$ follows because $\bar{\mathcal{E}}_{t}$ is a centered ellipsoid, and $(ii)$ from the total law of probability.
Using Lemma D.4, we conclude that with probability $1-\delta$, we have that
$\mathsf{SubOpt}(\hat{\pi})\leq\varepsilon(1+2\beta_{T+1}(\delta^{\prime}))+\frac{2\beta_{T+1}(\delta^{\prime})c(\delta^{\prime\prime})}{pT}\underbrace{\sum_{t=1}^{T}\mathbb{E}\left[{\left\lVert\phi(\tau_{t})-\phi(\tau_{t}^{\prime})\right\rVert_{V_{t}^{-1}}\mid\mathcal{F}_{t-1}}\right]}_{\leq C_{T}}$
$\leq$ $\left[{\sqrt{\kappa\left[{\log\left({\frac{5}{\delta}}\right)+d\log\left({\max\left\{{e,\frac{4eBLH_{\gamma}T}{d}}\right\}}\right)}\right]}+2\sqrt{\lambda}B}\right]\Bigg\{\dfrac{3}{\sqrt{T}}+\dfrac{2\sqrt{2d\log(10dT/\delta)}}{pT}$
$\cdot\left[{2\sqrt{T\left({2d\log\left(1\!+\!\frac{4TL^{2}H_{\gamma}^{2}}{d\lambda}\right)\!\!+\!\frac{12dL^{2}H_{\gamma}^{2}}{\log(2)\lambda}\log\left(1\!+\!\frac{4L^{2}H_{\gamma}^{2}}{\log(2)\lambda}\right)}\right)}+\frac{16LH_{\gamma}}{\sqrt{\lambda}}\log\left({\frac{10}{\delta}}\right)}\right]\Bigg\}$
Remark B.1 (Connection to G-optimal design).
Given a subset $\mathcal{X}\subset\mathbb{R}^{d}$, a G-optimal design selects $x_{1},\ldots,x_{T}\in\mathcal{X}$ to minimize $\max_{x\in\mathcal{X}}\left\lVert x\right\rVert_{V^{-1}}$ with $V=\sum_{t=1}^{T}x_{t}x_{t}^{\top}$. Under compactness and if $\operatorname{span}(\mathcal{X})=\mathbb{R}^{d}$, the Kiefer-Wolfowitz theorem [kiefer1960equivalence] yields the lower bound $\max_{x\in\mathcal{X}}\left\lVert x\right\rVert_{V^{-1}}\geq\sqrt{d/T}$, which is tight up to rounding. In the proof above, we show that Algorithm 2 guarantees $\max_{\pi,\pi^{\prime}}\left\lVert\phi(\pi)-\phi(\pi^{\prime})\right\rVert_{V_{T+1}^{-1}}=\mathcal{O}(d\log(dT/\delta)/\sqrt{T})$. Thus, ignoring the noise in $\phi(\tau)$ (handled via Freedman) and, for simplicity, the regularization $\lambda$, Algorithm 2 achieves a $\mathcal{O}(\sqrt{d}\log(dT/\delta))$-approximate G-optimal design with probability at least $1-\delta$.
@@
[…]
@@
Lemma C.1 (Number of design matrix updates).
@@
$\sum_{t=1}^{T}\mathbbm{1}\left[{V_{t+1}\neq V_{t}}\right]\leq\frac{d}{\log\left({1+C}\right)}\log\left({1+\frac{T(2LH_{\gamma})^{2}}{d\lambda}}\right).$